OpenStudy (anonymous):

solve by elimation. in a right triangle one angle is 54 degrees more than twice the other. find each acute angle.

Directrix (directrix):

Let x and (2x + 54) be the acute angles of the right triangle. The acute angles of a right triangle are complementary. 3x + 54 = 90 3x = 36 x = 12 2x + 54 = 2* 12 + 54= 78 The acute angles have measure 12 and 78.-->

OpenStudy (anonymous):

thanks

Directrix (directrix):

Glad to help.